OpenStudy (reverdeparadis):

I'm seriously stuck on this question. I've asked my friend and sister for help, but neither of them get it. Please answer ASAP. Thank you (: Table: x y 1 1 2 1 + 2 3 1 + 2 + 4 4 1 + 2 + 4 + 8 1.) Extend the table by writing expressions for y when x = 5, 6, and 7. 2.) Simplify each expression for y in your table from part 1 (above). What pattern do you notice in the simplified expressions? 3.) Write an equation that relates x and y. Use your equation to find the value of y when x = 15.

jimthompson5910 (jim_thompson5910):

when x = 5, the value of y will be 1+2+4+8+16 = 31 when x = 6, the value of y will be 1+2+4+8+16+32 = 63 I'm sure you're starting to see the pattern by now

jimthompson5910 (jim_thompson5910):

if not, then let me know

OpenStudy (reverdeparadis):

i understand the pattern in the expressions, but not in the sums..

jimthompson5910 (jim_thompson5910):

hint: compare these two sequences 1, 2, 4, 8, 16, 32, 64 1, 3, 7, 15, 31, 63

OpenStudy (reverdeparadis):

okay, so the first pattern is multiplying by 2, right? and i'm still not sure on the second one

jimthompson5910 (jim_thompson5910):

notice how we have 4 in the top row and 3 in the bottom row we have 8 in the top row, then 7 in the bottom we have 16 in the top row, then 15 in the bottom etc etc

OpenStudy (anonymous):

r u in connection academy

OpenStudy (anonymous):

when x = 5, the value of y will be 1+2+4+8+16 = 31 when x = 6, the value of y will be 1+2+4+8+16+32 = 63 I'm sure you're starting to see the pattern by now